OpenStudy (anonymous):

What is the solution to 3x^2+3x+5=0 (WILL GIVE MEDAL)

OpenStudy (shamil98):

Use the quadratic formula.

OpenStudy (anonymous):

I know how to do that i just dont know what to do after i put the numbers in

OpenStudy (shamil98):

\[x = \frac{ -b \pm \sqrt {b^2 - 4ac} }{ 2a }\] a = 3 , b = 3 , c = 5

OpenStudy (shamil98):

just plug them in and simplify.

OpenStudy (shamil98):

I'll get you started. \[x = \frac{ -3\pm \sqrt {3^2 - 4(3)(5)} }{ 2(3) }\]

OpenStudy (anonymous):

-3 +- sqrt(3^2 -4 (3)(5) -------------------- 2(3) ?? is that right?

OpenStudy (shamil98):

Yeah

OpenStudy (anonymous):

ok, im not sure how to simplify tho

OpenStudy (shamil98):

Simplifying. We get. \[x = \frac{ -3 \pm \sqrt{9-60} }{ 6 }\] \[x = \frac{ -3 \pm \sqrt{-51} }{ 6 }\] \[x = \frac{ -1 \pm \sqrt{-51} }{ 2 }\]

OpenStudy (anonymous):

ohh ok makes sense! Thank you so much!

OpenStudy (shamil98):

You're welcome.